Mikael Lindström , Department of Mathematics, Åbo Akademi University, FIN-20500, Åbo, Finland. E-mail address: mikael.lindstrom@abo.fi

Spectra of weighted composition operators on weighted

Banach spaces of analytic functions.

Joint work with Richard Aron

We determine the spectra of weighted composition operators acting on the weighted Banach spaces of analytic functions H_{v_{p}}^{∞} when the symbol ϕ has a fixed point in the open unit ball. As an application of this result we give the spectra of composition operators on Bloch type spaces.